# Misinformation is not about Bad Facts: An Analysis of the Production and Consumption of Fringe Content

### v1

@@ 5 Related Work
In this section, We investigate content-based misinformation detection methods and their associated limitations, as well as style-based approaches.
@@
Traditionally, misinformation detection techniques have relied on content-based information, such as encoded texts using language models. Alkhodair et al. (2020) proposed a recurrent neural network model for detecting rumors in breaking news using word embeddings. Similarly, Horne and Adali (2017) integrated three types of features, including stylistic elements, to classify fake news. Their findings revealed significant distinctions between fake and real news content. Meanwhile, Sarnovskỳ, Maslej-Krešňáková, and Ivancová (2022) concentrated on identifying fake news within the Slovak online sphere. They applied deep learning models to Word2Vec representations of the texts. While these models demonstrated significant performance, it is worth noting that the dataset predominantly comprised articles centered around COVID-19.
A significant shortcoming of content-based methods lies in the dynamic nature of misinformation topics. Models trained on predefined subjects may struggle to adapt to emerging themes. Furthermore, training deep learning models requires substantial data to mitigate the risk of overfitting. To address these limitations, Raza and Ding (2022) introduced a context-based model focusing on social aspects to identify fake news. This model incorporates users’ social interactions, such as comments on news articles, posts, and replies, as well as upvotes and downvotes. This approach could serve as a valuable complement, especially when engagement data is readily available.
@@
Whitehouse et al. (2022) showed that general-purpose content-based classifiers tend to overfit to specific text sources. They introduced a style-based classification approach; however, the proposed stylometric features leverage the categories in the General Inquirer (GI) dictionary, encompassing content-specific words related to religion and politics. We omit such specific categories in this work when employing the Linguistic Inquiry and Word Count (LIWC).
More recently, Kumarage et al. (2023) demonstrated the effective detection of AI-generated texts in Twitter timelines using various stylistic signals. Their study showed that classifiers employing the proposed stylometric features outperformed Bag of Words and Word2Vec embeddings. Notably, they mentioned that among the stylometric features, punctuation and phraseology features proved to be the most significant. While these findings are motivating, their research primarily focuses on distinguishing between human and AI-authored content within a given Twitter timeline. In contrast, our goal is to discriminate between different writing styles, such as intimate and expert styles. In another attempt to verify authorship using style representations, Wegmann, Schraagen, and Nguyen (2022), content-controlled style representations were proposed. Authors demonstrated that by controlling different levels of contents, e.g., authorship verification within texts from the same conversation or the same domain, the performance varies. They also manually inspected the resulting clusters to find out what styles are learned such as ‘punctuation’. In our proposed work, we attempt to learn and distinguish styles used by extreme groups since we observed that certain styles were strategically adopted by these groups to reach vulnerable demographics in online misinformation space.
@@ 6 Conclusion and Discussion
In this paper, we showed that misinformation is conveyed through styled messages by detecting styles in extreme online communities and being able to distinguish the communities using styles rather than content. Specifically, Facebook pages which share misinformation can be distinguished by the style of the posts made within each group. The classification results showed that stylistic measures can outperform the content-based classifier. This is intriguing since the two groups, especially an anti-vaccination group, can be easily identified through the analysis of content-related vocabulary.
We also evidenced that far-right produced contents are significantly different from what far-right online users consume. We showed that the far-right production and the far-right consumption can be distinguished by using stylistic features. This indicates that misinformation consumers prefer certain styles of information as opposed to the contents of information.
This study has limitations across multiple dimensions. Firstly, the Facebook group dataset is constrained in its post volume. Additionally, the availability of labeled style data was even more scarce, limiting the generalizability of the findings. Our plan in the future is to apply active learning process to label more posts efficiently. Secondly, we focused only on styles of messages in order to distinguish and observe effects styled text. While studying online misinformation space, we noticed that misinformation producers not only strategically adopt styles to effectively spread the information, but also utilise formats such as graphics, videos and reports. By considering these facets of misinformation packaging, we can enhance the results further. Lastly, we intend to explore and develop our understanding of the vulnerable demographics in online spaces. We observed that extremist online groups aimed at these personas tailor their content according to these preferences. Better understanding the patterns and styles of misinformation that the vulnerable demographics are attracted to can serve as a guiding reference for policy makers, suggesting the adoption of a comparable strategy to effectively counter the dissemination of misinformation.
@@
All data that we obtained was publicly available at the time of data collection. We discarded deleted, protected, and private content at the time of analysis. Therefore, the analyses reported in this work does not compromise any user privacy.
@@
By exploring landscape of Australian news media and online misinformation space, we hope to raise awareness of how the language styles used in these online communities may appeal to vulnerable demographics. Also, understanding how misinformation is styled and packaged showed insights on how we can approach affected population with reliable information. Awareness goes both ways, though, and the results presented in the paper could also lead to malicious users learning about ways to easily manipulate the target audience. However, based on our observation, creators of misinformation with intention to reach a target population are already styling their messages accordingly. By showing that misinformation consumers are attracted to styled messages, we hope any entities involved in misinformation intervention campaigns utilize these findings.
@@ Appendix A Details of Stylistic measures
We summarize the linguistic measures we used in Table 9. The full list of features of StyloMetrix is found here [^5] https://github.com/ZILiAT-NASK/StyloMetrix/blob/v0.1.0/resources/metrics˙list˙en.md.
@@
Summary
LIWC is a transparent text analysis program that counts words in psychologically meaningful categories
Grievance dictionary assess grievance-fuelled communications through language
StyloMetrix is a tool for creating text representations as StyloMetrix vectors
# features
89
22
175
features
Segment, WC, Analytic, Clout, Authentic, Tone, WPS, BigWords, Dic, Linguistic, function, pronoun, ppron, i, we, you, shehe, they, ipron, det, article, number, prep, auxverb, adverb, conj, negate, verb, adj, quantity, Drives, affiliation, achieve, power, Cognition, allnone, cogproc, insight, cause, discrep, tentat, certitude, differ, memory, Affect, tone_pos, tone_neg, emotion, emo_pos, emo_neg, emo_anx, emo_anger, emo_sad, swear, Social, socbehav, prosocial, polite, conflict, moral, comm, socrefs, substances, risk, curiosity, allure, Perception, attention, motion, space, visual, auditory, feeling, time, focuspast, focuspresent, focusfuture, Conversation, netspeak, assent, nonflu, filler, AllPunc, Period, Comma, QMark, Exclam, Apostro, OtherP
deadline, desperation, fixation, frustration, god, grievance, hate, help, honour, impostor, jealousy, loneliness, murder, paranoia, planning, relationship, soldier, suicide, surveillance, threat, violence, weaponry

### v2

@@ 5 Related Work
In this section, we investigate content-based misinformation detection methods, their associated limitations, and style-based approaches.
@@
Traditionally, misinformation detection techniques have relied on content-based information, such as encoded texts using language models. Alkhodair et al. (2020) proposed a recurrent neural network model for detecting rumors by utilizing Word2Vec (Mikolov et al. 2013) representation. They showed that their model outperforms the state-of-the-art sequential classifier from Zubiaga, Liakata, and Procter (2016). They then applied their model to emerging breaking news in a real-time Twitter stream. The F1 scores for the two case studies were $0.757$ and $0.791$. Our style-based classifier achieved comparable results when distinguishing far-right production and far-right consumption. Similarly, Horne and Adali (2017) integrated three types of features, including stylistic elements, to classify fake news. Their findings revealed significant distinctions between fake and real news content. Meanwhile, Sarnovskỳ, Maslej-Krešňáková, and Ivancová (2022) concentrated on identifying fake news within the Slovak online sphere. They applied deep learning models to Word2Vec representations of the texts. While the models demonstrated remarkable performance (best model reaching the accuracy of $98.93\%$), it is worth noting that the dataset predominantly comprised articles centered around COVID-19.
A shortcoming of content-based methods lies in the dynamic nature of misinformation topics. Models trained on predefined subjects may struggle to adapt to emerging themes. Furthermore, training deep-learning models requires substantial data to mitigate the risk of over-fitting. To address these limitations, Raza and Ding (2022) introduced a context-based model focusing on social aspects to identify fake news. This model incorporates users’ social interactions, such as comments on news articles, posts, and replies, as well as upvotes and downvotes. This approach can serve as a valuable complement, especially when engagement data is readily available.
@@
Whitehouse et al. (2022) showed that general-purpose content-based classifiers tend to overfit to specific text sources. In response, the authors proposed a style-based classification approach. The proposed stylometric features, however, leverage the categories in the General Inquirer (GI) dictionary, encompassing content-specific words related to religion and politics. To factor out content as much as possible, we omit such specific categories when employing the Linguistic Inquiry and Word Count (LIWC).
More recently, Kumarage et al. (2023) demonstrated the effective detection of AI-generated texts in Twitter timelines using various stylistic signals. Their study showed that classifiers employing the proposed stylometric features outperformed Bag of Words (Manning, Raghavan, and Schütze 2008) and Word2Vec embeddings (Mikolov et al. 2013). Notably, they mentioned that among the stylometric features, punctuation and phraseology features proved to be the most significant. While these findings are motivating, their research primarily focuses on distinguishing between human and AI-authored content within a given Twitter timeline. In contrast, we aim to discriminate between different writing styles, such as intimate and expert styles.
In another attempt to verify authorship using style representations content-controlled style representations have been proposed (Wegmann, Schraagen, and Nguyen 2022). The authors demonstrated that performance varies when controlling for different levels of contents, e.g., authorship verification within texts from the same conversation or the same domain. They used a clustering algorithm on text samples and manually inspected the resulting clusters of texts to find out what styles were learned, such as ‘punctuation’. In our proposed work, we attempt to learn and distinguish styles used by extreme groups since we observed that these groups strategically adopted certain styles to reach vulnerable demographics in online misinformation space. This is different from learning innate styles of individuals which maybe more subtle.
@@ 6 Conclusion and Discussion
We have shown that misinformation is conveyed through styled messages by detecting styles in extreme online communities and being able to distinguish the communities using styles rather than content. Specifically, Facebook pages that share misinformation can be distinguished by the style of the posts made within each group. The classification results showed that stylistic measures can outperform the content-based classifier. This is intriguing since the two groups, especially an anti-vaccination group, can be easily identified by analyzing content-related vocabulary.
We also evidenced that content produced by the far-right differs significantly from far-right online users’ consumption. This is to say that while producers may cover a wide range of topics, users selected a narrow set of articles to share, which had consistent style features. Thus, we showed that the far-right production and the far-right consumption can be distinguished by using stylistic features. This indicates that misinformation consumers prefer certain styles of information as opposed to the contents of information.
This study has limitations across multiple dimensions. Firstly, the Facebook group dataset is constrained in its post volume. Additionally, the availability of labeled style data was even more scarce, limiting the generalizability of the findings. Our plan in the future is to apply an active learning process to label more posts efficiently. Secondly, we focused only on the styles of messages in order to distinguish and observe the effects of styled text. While studying online misinformation, we noticed that misinformation producers strategically adopt styles to effectively spread the information and utilize formats such as graphics, videos and reports. We can enhance the results by considering these facets of misinformation packaging.
Lastly, we intend to explore and develop our understanding of the vulnerable demographics in online spaces. We observed that extremist online groups aimed at these personas tailor their content to these preferences. A better understanding of the patterns and styles of misinformation the vulnerable demographics are attracted to can serve as a guiding reference for policymakers, suggesting adopting a comparable strategy to effectively counter misinformation dissemination.
@@
All data that we obtained was publicly available at the time of data collection. We discarded deleted, protected, and redacted content at the time of analysis. Therefore, the analyses reported in this work does not compromise any user privacy. This project was approved by the Human Ethics Committee of our institution (approval number redacted for review).
@@
By exploring the landscape of Australian news media and the online misinformation space, we hope to raise awareness of how language styles used in online communities may appeal to vulnerable demographics. Understanding how misinformation is styled and packaged provides insights into how we can approach affected populations with reliable information. Awareness goes both ways, however, and the results presented in the paper could also lead to malicious users learning ways to easily manipulate a target audience. Based on our observations, however, creators of misinformation with the intention to reach a target population are already styling their messages accordingly. By showing that misinformation consumers are attracted to styled messages, we hope any entities involved in misinformation intervention campaigns can utilize these findings.
@@ Appendix A Details of Stylistic measures
We summarize the linguistic measures and list the features that we used in Section 4. The full list of features of StyloMetrix is found here [^7] https://github.com/ZILiAT-NASK/StyloMetrix/blob/v0.1.0/resources/metrics˙list˙en.md.
@@
Summary LIWC is a transparent text analysis program that counts words in psychologically meaningful categories Grievance dictionary assess grievance-fuelled communications through language StyloMetrix is a tool for creating text representations as StyloMetrix vectors
# features 89 22 175
features Segment, WC, Analytic, Clout, Authentic, Tone, WPS, BigWords, Dic, Linguistic, function, pronoun, ppron, i, we, you, shehe, they, ipron, det, article, number, prep, auxverb, adverb, conj, negate, verb, adj, quantity, Drives, affiliation, achieve, power, Cognition, allnone, cogproc, insight, cause, discrep, tentat, certitude, differ, memory, Affect, tone_pos, tone_neg, emotion, emo_pos, emo_neg, emo_anx, emo_anger, emo_sad, swear, Social, socbehav, prosocial, polite, conflict, moral, comm, socrefs, substances, risk, curiosity, allure, Perception, attention, motion, space, visual, auditory, feeling, time, focuspast, focuspresent, focusfuture, Conversation, netspeak, assent, nonflu, filler, AllPunc, Period, Comma, QMark, Exclam, Apostro, OtherP deadline, desperation, fixation, frustration, god, grievance, hate, help, honour, impostor, jealousy, loneliness, murder, paranoia, planning, relationship, soldier, suicide, surveillance, threat, violence, weaponry